OpenStudy (anonymous):

What are the zeroes of f(x) = x2 - x - 2? A) x = -2, 1 B) x = 2, -1 C) x = -2, -1 D) x = 2, 1

OpenStudy (anonymous):

Is it \[x ^{2}-x-2?\] Because if thats what you're asking the answer is 2, -1

OpenStudy (anonymous):

Yes it is. Thank you.

OpenStudy (anonymous):

Yeah you just have to factor it f(x)=(x-2)(x+1) f(x) is 0 when x =2,-1